# COMMA: A Communicative Multimodal Multi-Agent Benchmark

### v3

@@ 4.1 Experimental Setup
In this section, we describe the experimental settings of our multi-agent interaction environment where two distinct agents, namely the Solver agent and the Expert agent, engage in iterative dialogue sessions. The primary aim of this setup is to assess the collaborative problem-solving capabilities between different agents. During our experiments, we limit the number of conversation turns to 10 and the number of mistakes to 3, allowing for a unified and systematic assessment of interactions. The puzzle set used in evaluation consists of 100 fixed but different initializations of each of the 10 puzzles, resulting in 1000 total conversations. For most models, we use greedy decoding when available to maintain consistent agent output across different runs of the same puzzle. However, for reasoning models we set the temperature to 0.6 to avoid endless repetition. All inference is run on a single NVIDIA A100 GPU with 80GB VRAM. We parse the solver’s chosen actions at each conversation turn using exact string matching and directly perform the action on the interface if the solver outputs a valid action. Our exact prompts for both the solver and expert agent can be found in Appendix D.
@@ 4.2 Evaluation Metrics
  • Partial Success Rate (PSR): Because our benchmark includes puzzles with multi-step reasoning, some puzzles can have a more precise success rate evaluation. For these multi-step puzzles, we assign the solver a number between 0 and 100 to indicate its progress towards the solution, and average this number across puzzles to obtain a partial success rate. For single-step puzzles, the partial success rate is identical to the success rate.
@@ 5.3 Fine-grained Analysis
Figure 5: Illustration of the tradeoff between performance and token usage on our benchmark. Agents in the top left corner have a better balance between performance and token usage. Among the AI agents, GPT-4o has the best balance, but still falls far below human-level efficiency.
@@
Setting Average Partial Success Rate % (↑) Wire Telehealth Who LED Memory Keypad Password Color Maze Atm Overall Full Memory 98 ± 1.4 64 ± 4.7 72 ± 4.5 32 ± 3.4 31 ± 3.9 27 ± 2.5 2 ± 1.4 10 ± 1.6 33 ± 3.6 47 ± 5.0 41.74 ± 1.4 No Memory 95 ± 2.2 56 ± 4.8 71 ± 4.6 37 ± 3.9 24 ± 3.1 24 ± 2.3 7 ± 2.6 8 ± 1.3 3 ± 1.7 45 ± 5.5 36.20 ± 1.4 Single Agent 82 ± 3.9 74 ± 4.3 56 ± 5.0 66 ± 3.6 24 ± 3.0 34 ± 3.7 11 ± 3.1 17 ± 1.6 33 ± 5.0 8 ± 2.7 40.41 ± 1.4
Table 2: Impact of changing information access on performance in our benchmark for a GPT-4o agent. Interestingly, providing the agent with full information slightly worsens the overall performance. However, performance on more reasoning-heavy tasks improves. Withholding memory from past conversations hurts performance across most puzzles.
Effect of Memory and View Asymmetry
To isolate the effects of memory input and view asymmetry, we run two ablations on the collaborative setting. First, we remove episodic memory by restricting the agent to only the latest conversation message. Second, we remove view asymmetry by giving a single agent full access to both the solution manual and the image. Results are shown in Table 2. Removing memory significantly reduces performance, especially on tasks requiring reasoning across multiple turns (e.g., maze and memory puzzles). This shows that collective memory is crucial for recalling information over interactions; without it, the agent repeats itself, contradicts earlier steps, or loses progress. Interestingly, the single-agent setting with full information also performed slightly worse overall, though it sometimes outperformed collaboration on reasoning-heavy tasks (e.g., LED, Color). The large performance gap across tasks suggests that while some benefit from collaboration, others are better handled by individual reasoning.
@@ Acknowledgement
Research reported in this publication was partially supported by the National Institute Of Biomedical Imaging And Bioengineering of the National Institutes of Health under Award Number R01EB033782. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health.

### v4

@@ 4.1 Experimental Setup
In this section, we describe the experimental settings of our multi-agent interaction environment where two distinct agents, namely the Solver agent and the Expert agent, engage in iterative dialogue sessions. The primary aim of this setup is to assess the collaborative problem-solving capabilities between different agents. During our experiments, we limit the number of conversation turns to 10 and the number of mistakes to 3, allowing for a unified and systematic assessment of interactions. The puzzle set used in evaluation consists of 100 fixed but different initializations of each of the 10 puzzles, resulting in 1000 total conversations. For most models, we use greedy decoding when available to maintain consistent agent output across different runs of the same puzzle. However, for reasoning models we set the temperature to 0.6 to avoid endless repetition. All inference is run on a single NVIDIA A100 GPU with 80GB VRAM. We parse the solver’s chosen actions at each conversation turn using exact string matching and directly perform the action on the interface if the solver outputs a valid action. Exact prompts for both agents are in Appendix D.
@@ 4.2 Evaluation Metrics
  • Partial Success Rate (PSR): Because our benchmark includes puzzles with multi-step reasoning, some puzzles can have a more precise success rate evaluation. For these multi-step puzzles, we assign the solver a number between 0 and 100 to indicate its progress towards the solution, and average across puzzles for a partial success rate. For single-step puzzles, the partial success rate equals the success rate.
@@ 5.3 Fine-grained Analysis
Figure 5: Tradeoff between performance and token usage on our benchmark. Agents in the top left corner achieve a better balance between performance and efficiency. Among AI agents, GPT-4o performs best, but still far below human efficiency.
@@
Setting Average Partial Success Rate % (↑) Wire Telehealth Who LED Memory Keypad Password Color Maze Atm Overall Full Memory 98 ± 1.4 64 ± 4.7 72 ± 4.5 32 ± 3.4 31 ± 3.9 27 ± 2.5 2 ± 1.4 10 ± 1.6 33 ± 3.6 47 ± 5.0 41.74 ± 1.4 No Memory 95 ± 2.2 56 ± 4.8 71 ± 4.6 37 ± 3.9 24 ± 3.1 24 ± 2.3 7 ± 2.6 8 ± 1.3 3 ± 1.7 45 ± 5.5 36.20 ± 1.4
Table 2: Impact of changing information access on performance in our benchmark for a GPT-4o agent. Withholding memory from past conversations hurts performance across most puzzles, causing a significant overall performance drop.
Effect of Conversation Memory
To isolate the effects of memory input, we performed an experiment in the collaborative setting by removing episodic memory. Specifically, we restrict each agent to only have access to the latest conversation message. The results are shown in Table 2. Removing memory significantly reduces performance, especially on tasks requiring reasoning across multiple turns (e.g., maze and memory puzzles). This shows that collective memory is crucial for recalling information over interactions; without it, the agent repeats itself, contradicts earlier steps, or loses progress.
@@ Acknowledgement
Research reported in this publication was partially supported by the National Institute Of Biomedical Imaging And Bioengineering of the National Institutes of Health under Award Number R01EB033782, and the National Science Foundation under Award Number IIS-2449768. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health and the National Science Foundation.